These problems make use of the following data for hypothetical economy

C=140+6(Y-T) I=200 I has the ( -) symbol on top G=150 T=150

All variables are in units of constant dollars per year in units of constant dollars per year,which we will represent by the symbol $ symbol. unless the problem states otherwise ,always go back to these original data calculate the equilibrium level of output Derive the equation for aggregate demand. Calculate the level of demand and unplanned inventory change when output is$500 per year and describe the supply response. Repeat when the output is $1200 per year. The government wants to increase GDP by $250. calculate the level of government spending required.
Answer #1

Y=140+0.6(y-150)+200+150

Y=140+0.6y-90+200+150=400+0.6y

Y=400/0.4=1000

AD=C+I+G

I=200( fixed)

G=150

AD=140+0.6(y-150)+200+150

AD=400+0.6Y

At y=500

AD=400+0.6*500=400+300=700(aggregate demand)

Unplanned inventory change=equilibrium Y - AD=1000-700=300

Y=1200

AD=400+0.6*1200=400+720=1120

Unplanned inventory change=1000-1120=-120

Government spending Multiplier=1/(1-mpc)=1/(1-0.6)=1/0.4=2.5

∆Y=250

∆G=∆y/government spending multiplier=250/2.5=100

New G=150+100=250